# Principles of hydrodynamic particle manipulation in internal Stokes flow

### v1

@@ 1 Introduction
Microfluidic particle manipulation aims at transportation, separation, trapping, and enrichment (Sajeesh & Sen, 2014; Lu et al., 2017) and can be achieved through various approaches. Many techniques exploit certain particles’ response to external forces, e.g. electrical (Xuan, 2019), optical (Lenshof & Laurell, 2010), and magnetic techniques (Van Reenen et al., 2014). However, not all particles of interest will be susceptible to these, which is why there is a continued interest in manipulation based solely on hydrodynamic forces (Karimi et al., 2013). Most notably, techniques that use particle inertia have gained prominence (Di Carlo et al., 2007; Di Carlo, 2009; Agarwal et al., 2018) and quantitative theories have been developed beyond classical equations of motion (Maxey & Riley, 1983) to rigorously describe the effect of inertial forces in both the background flow and the disturbance flow around the particle (Agarwal et al., 2021; Agarwal, 2021). Acoustofluidics (Bruus et al., 2011; Laurell et al., 2007; Friend & Yeo, 2011), a subset of inertial effects, has also been quantitatively connected recently (Agarwal et al., 2024). Despite the description of such forces in simple flow fields that is now known analytically, many practical cases still lack a fundamental quantitative theory on how devices based on hydrodynamic effects work.
@@ 2 HYDRODYNAMIC MODEL OF PARTICLE-WALL INTERACTION IN STOKES FLOW
Generally, studies of force-free particles result in predictions for the deviation of the particle velocity from the background fluid velocity it is embedded in, i.e., the non-passive part of the particle motion. This velocity correction can be decomposed into effects in the direction parallel to the boundary and in the perpendicular direction. A comprehensive analysis of the equation of motion of a force-free particle with wall-normal velocity corrections is given by Rallabandi et al. (2017). Other work has described wall-parallel velocity corrections far from and near the wall (Ekiel-Jeżewska & Wajnryb, 2006; Pasol et al., 2011). Here, we make use of or recompute velocity corrections in both the wall-parallel and wall-perpendicular directions, constructing uniformly valid expressions for all distances from the wall in order to compute particle trajectories.
@@ 2.1 Moffatt Eddies
The Moffat parallel plate solution consists of a series of alternating congruent vortices that take up the height of the channel (Fig. 2(b) shows streamline contours) and whose strength decays exponentially with distance from a stirrer on the far left. The corresponding stream function $\psi$ is known analytically (asymptotically far from the stirrer) and has the form
@@ 2.3 Wall-parallel corrections to the particle velocity
Figure 3: (a) Moffatt eddy flow with pairs of counterrotating vortices taking up the channel height, from the antisymmetric stream function (2). (b) Wall-parallel flow modification factor $f(\Delta)$ at large $\Delta$. (c) Plot of $f(\Delta)$ at small $\Delta$. (d) Example of normal particle velocity as a function of $\Delta$, for $x=1$ and $a_{p}=0.1$: far from the wall, the model follows the particle-expansion velocity $v_{p}^{PE}$ from (10), while for $\Delta<0.5$ (inset) the variable-expansion approach of (12) is used.
@@ 2.4 Wall-normal corrections to the particle velocity
For our case of a force-free particle, we set $\boldsymbol{F_{\perp}}=0$ in (5) as well as $\boldsymbol{e}_{\perp}=\pm\boldsymbol{e}_{y}$ (for the wall at $y=\mp 1$, repectively). The resulting equation can be solved for the wall-normal particle velocity $v_{p}^{PE}$ to isolate the wall-normal velocity corrections $W_{y}^{\pm}$ due to the presence of walls at $y=\pm 1$, i.e.,
@@
and it is understood that $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ are evaluated at arguments $\Delta$ from (7) defined by $h=1\pm y_{p}$ for the walls at $y=\mp 1$, respectively. Note that the last term of (11) explicitly subtracts the Faxen correction so that each $W_{y}$ term vanishes as $\Delta\to\infty$. When the particle approaches a wall, we anticipate approaches to $\Delta\ll 1$. In the wall-parallel direction, we still use the expression for $W_{x}$ from equation (6), which includes the lubrication limit at very small $\Delta$. However, in the wall-normal direction, the formalism relying on Taylor expansion of the flow around the particle center $v_{p}^{PE}$ is not accurate enough to describe the particle motion – this is easily seen because the predicted particle normal velocity from (10) does not vanish when the particle touches the wall. In (Rallabandi et al., 2017), it was shown that for $\Delta\ll 1$, the no-penetration boundary condition can be enforced by replacing the background flow field in (11) by its quadratic expansion around the point on the wall closest to the particle. However, we find that this wall-expansion formalism will not smoothly transition to the expression (11) when $\Delta\sim 1$.
Therefore, we generalize and improve on the wall-expansion procedure here and construct second-order expansions of $v(x,y)$ around variable expansion points $x_{E}=x_{p},y_{E}=y_{E}(y_{p})$, i.e.,
@@
We omit the linear in $x$ and mixed terms because these will never give nonzero contributions when evaluated within our formalism. The expansion point must coincide with the nearest point on the wall when the particle is touching, and with the particle $y$-position as $\Delta\to 1$ to consistently merge into the particle-expansion formalism. Thus, for a wall at $y=-1$ and a particle at $y_{p}$ we set
@@
Note that the effects of the wall at $y=+1$ are negligible. The wall correction is now
@@
Using equations (6), (8), and (10) – (16) in the dynamical system (5), We have thus established a formalism for computing particle trajectories in the presence of channel wall effects for arbitrary Stokes background flow.
@@ 3.2.1 Linear stability analysis
where the constant $c$ can vary depending on the specific flow, but is ${\cal O}(1)$ (Goldman et al., 1967a; Ghalia et al., 2016), cf. Sec. 2.3. Near the fixed point of the Faxen field, the $\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}$ term in $W_{y}$ dominates the others ($\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}\propto\mathcal{O}(a_{p}^{3}),\frac{\mathcal{C}_{large}}{\mathcal{A}_{large}}\propto\mathcal{O}(a_{p}^{5}),\frac{\mathcal{D}_{large}}{\mathcal{A}_{large}}\propto\mathcal{O}(a_{p}^{4})$) and is the only one contributing to $\mathcal{O}(a_{p}^{3})$ in (18). This term is still of higher $a_{p}$ order than the Faxen correction so that, in the limit of large $\Delta$ and small $a_{p}$, the wall correction is perturbatively small.
@@ 3.3 Particle motion and manipulation in a clockwise vortex
However, the ability to concentrate particles on a limit cycle trajectory without inertial effects purely because of the background flow geometry is of fundamental interest. It is encouraging that the location of this limit cycle for small particles can be determined entirely within the large $\Delta$ approximation, i.e., without the intricate details of near-wall corrections or lubrication limits. This gives confidence in not only the qualitative but also the quantitative description of the phenomenon: when placed in certain bounded vortical Stokes flows, small spherical particles, even when neutrally buoyant, will eventually accumulate on well-defined closed trajectories.
@@ 3.4 Particle motion and manipulation in a counterclockwise vortex
Figure 6(b) exemplifies a particle trajectory that approaches very close to the wall, with a nearly normal initial approach and a subsequent nearly parallel piece of trajectory. The semi-logarithmic plot vs. time in Fig. 6(c) shows that the ”near-parallel” portion contains a stretch of approximately exponential decay of the (relative) gap $\Delta$. Indeed, this behavior follows from a series to leading order in small $\Delta$ of (15), i.e., the wall-expansion limit of $y_{E}\to-1$. Here, (14) simplifies to
@@ 4 Conclusions
We have quantitatively shown that force-free particles placed in an internal Stokes flow can be systematically displaced in a variety of ways through purely hydrodynamic interactions with the enclosing walls. As a particle cannot contact a wall in finite time, a typical trajectory has a portion of approaching the wall and a portion of receding, governed by the background flow and the wall corrections derived from it. If the approaching and receding parts of the flow are symmetric, displacement effects off a streamline will cancel out. Thus, a Stokes-flow-inducing net particle displacement must break wall-parallel symmetry.
For practical use, vortical flows are advantageous, where the effects of many approaching/receding events can accumulate. However, if the vortex is symmetrically confined between walls, the net displacement effects induced by both walls will again cancel, and the particle trajectory must close. For net cumulative displacement, the vortex must, therefore, break wall-normal symmetry as well, for example, by being confined between no-slip and no-stress boundaries. A net-transport internal Stokes flow thus breaks symmetries in both directions by particle-wall interactions, which act as perturbations on the Faxen flow trajectories. This is in agreement with recent experimental results that record permanent displacements of fibers passing by obstacles of symmetry-broken shape (Li et al., 2024). For such non-spherical particles (whether rigid or elastic), the effect of a single particle-boundary encounter could be much enhanced, as the wall effects affect different parts of the particle differently, and the overall torque balance will influence the resulting displacement. The study of this case will be the subject of future work.
@@
Let us estimate the rate of particle drift across streamlines in dimensional terms. Taking the length scale (channel half-width) as $H=50\mu$m, we want to ensure that the channel Reynolds number $Re=HU/\nu$ for aqueous solutions ($\nu\sim 10^{-6}$m2/s) is less than $10^{-2}$ along the particle trajectories, even in regions of the highest speed $U$. Given this constraint and assuming $a_{p}=5\mu$m, the time scale $\tau$ from (24) is $\approx 57$s, so that a significant radial displacement is expected on a time scale of minutes. For larger particles, this time scale decreases rapidly ($\propto a_{p}^{-3}$), and the process can also be sped up by using a solution of higher viscosity.
The Moffatt eddy flow discussed here is convenient because of its closed analytical form. However, it would be extremely difficult to set up experimentally due to the exponential decay of the flow speed with $x$ and the need to drive the flow far away from the field of view. The exponential dependence of speed on coordinates also makes the Reynolds number constraint for Stokes flow extremely stringent. Nevertheless, the principles of displacement and the scaling of the wall effects are not specific to this flow and will be robust in any symmetry-breaking vortical Stokes flow.
@@
It is instructive to compare the scaling of the migration velocity of small particles far from the wall ($\Delta\gg 1$) with that of approaches using particle inertia. Various approaches quantifying the effect of steady inertia (Ho & Leal, 1974; Schonberg & Hinch, 1989; Di Carlo et al., 2007; Hood et al., 2015) agree that for force-free particles, the migration speed scales with $a_{p}^{3}$, which agrees with our result (18). In approaches using oscillatory inertia (Agarwal et al., 2021, 2024), the scaling varies from $a_{p}^{3}$ for small Stokes number to $a_{p}^{4}$ for large Stokes number. Thus, the strictly non-inertial effects described in the present work show scaling equal to or even more favorable than inertial techniques for small particles.

### v2

@@ 1 Introduction
Microfluidic particle manipulation aims at transportation, separation, trapping, and enrichment (Sajeesh & Sen, 2014; Lu et al., 2017) and can be achieved through various approaches. Many techniques exploit certain particles’ response to external forces, e.g. electrical (Xuan, 2019), optical (Lenshof & Laurell, 2010), and magnetic techniques (Van Reenen et al., 2014). However, not all particles of interest will be susceptible to these, which is why there is a continued interest in manipulation based solely on hydrodynamic forces (Karimi et al., 2013). Most notably, techniques that use particle inertia have gained prominence (Di Carlo et al., 2007; Di Carlo, 2009; Agarwal et al., 2018) and quantitative theories have been developed beyond classical equations of motion (Maxey & Riley, 1983) to rigorously describe the effect of inertial forces in both the background flow and the disturbance flow around the particle (Agarwal et al., 2021; Agarwal, 2021). Recent work by Agarwal et al. (2024) also integrates the important case of acoustofluidic particle manipulation (Bruus et al., 2011; Laurell et al., 2007; Friend & Yeo, 2011) as a particular limit of inertial particle manipulation. Despite the description of such forces in simple flow fields that is now known analytically, many practical cases still lack a fundamental quantitative theory on how devices based on hydrodynamic effects work.
@@ 2 HYDRODYNAMIC MODEL OF PARTICLE-WALL INTERACTION IN STOKES FLOW
Generally, studies of force-free particles result in predictions for the deviation of the particle velocity from the background fluid velocity it is embedded in, i.e., the non-passive part of the particle motion. This velocity correction can be decomposed into effects in the direction parallel to the boundary and in the perpendicular direction. A comprehensive analysis of the equation of motion of a force-free particle with wall-normal velocity corrections is given by Rallabandi et al. (2017). Other work has described wall-parallel velocity corrections far from and near the wall (Ekiel-Jeżewska & Wajnryb, 2006; Pasol et al., 2011). Here, we derive an original formalism for the displacement of spherical particles in Stokes flows, expanding on existing work to arrive at an equation of motion applicable for all particle-wall distances. We will demonstrate several modes of systematic (net) particle displacement across streamlines due to wall interaction effects, a phenomenon not previously acknowledged widely. While particles can never cross streamlines in unidirectional Stokes flow (Bretherton, 1962), streamline crossing should generally be expected in the presence of wall-normal flow components. A simple example is a particle very close to a wall transported in a channel flow undergoing contraction – the particle cannot stay on its initial streamline without penetrating the wall.
@@ 2.1 Moffatt Eddies
The Moffatt parallel plate solution consists of a series of alternating congruent vortices that take up the height of the channel (Fig. 2(b) shows streamline contours) and whose strength decays exponentially with distance from a stirrer on the far left. The corresponding stream function $\psi$ is known analytically (asymptotically far from the stirrer) and has the form
@@ 2.3 Wall-parallel corrections to the particle velocity
Figure 3: (a) Moffatt eddy flow with pairs of counterrotating vortices taking up the channel height, from the antisymmetric stream function (2). (b) Wall-parallel flow modification factor $f(\Delta)$ at large $\Delta$. (c) Plot of $f(\Delta)$ at small $\Delta$. (d) Example of normal particle velocity as a function of $\Delta$, for $x=1$ and $a_{p}=0.1$ in the flow of (a): far from the wall, the model follows the particle-expansion velocity $v_{p}^{PE}$ from (10), while for $\Delta<0.5$ (inset) the variable-expansion approach of (12) is used.
@@ 2.4 Wall-normal corrections to the particle velocity
For our case of a force-free particle, we set $\boldsymbol{F_{\perp}}=0$ in (5) as well as $\boldsymbol{e}_{\perp}=\pm\boldsymbol{e}_{y}$ (for the wall at $y=\mp 1$, repectively). The resulting equation can be solved for the wall-normal particle velocity $v_{p}^{PE}$ – the superscript PE stands for particle expansion as we expand the background flow around ${\bf x}_{p}(t)$. Writing the wall-normal velocity corrections $W_{y}^{\pm}$ due to the presence of both walls at $y=\pm 1$ separately, we have
@@
and it is understood that $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ are evaluated at arguments $\Delta$ from (7) defined by $h=1\pm y_{p}$ for the walls at $y=\mp 1$, respectively. Note that the last term of (11) explicitly subtracts the Faxen correction so that each $W_{y}$ term vanishes as $\Delta\to\infty$. When the particle approaches a wall closely, we enter the regime of $\Delta\ll 1$. In the wall-parallel direction, we still use the expression for $W_{x}$ from equation (6), which includes the lubrication limit at very small $\Delta$. However, in the wall-normal direction, the formalism relying on Taylor expansion of the flow around the particle center $v_{p}^{PE}$ is not accurate enough to describe the particle motion – this is easily seen because the predicted particle normal velocity from (10) does not vanish when the particle touches the wall. In (Rallabandi et al., 2017), it was shown that for $\Delta\ll 1$, the no-penetration boundary condition can be enforced by replacing the background flow field in (11) by its quadratic expansion around the point on the wall closest to the particle. However, we find that this wall-expansion formalism will not smoothly transition to the expression (11) when $\Delta\sim 1$, because the exponential dependence of the flow field (2) with substantial $|k|$ compromises the accuracy of the quadratic expansion even at relatively short distances from the wall.
Therefore, we generalize and improve the transition to $v_{p}^{PE}$ at small $\Delta$ by constructing the quantity $v^{VE}(x,y)$, the second-order expansion of $v(x,y)$ around variable expansion points $x_{E}=x_{p},y_{E}=y_{E}(y_{p})$, i.e.,
@@
We omit the linear in $x$ and mixed terms because these never give nonzero contributions when evaluated within our formalism. The expansion point must coincide with the nearest point on the wall when the particle is touching, and with the particle $y$-position as $\Delta\to 1$ to consistently merge into the particle-expansion formalism. Thus, for a wall at $y=-1$ and a particle at $y_{p}$ we set
@@
Note that the effects of the wall at $y=+1$ are negligible here. The wall correction is now
@@
Using equations (6), (8), and (10) – (16) in the dynamical system (5), We have thus established a formalism for computing particle trajectories in the presence of channel wall effects for arbitrary Stokes background flow. To the authors’ knowledge, the present work is the first to formulate a closed hydrodynamics-based equation of motion for particles entrained in wall-bounded Stokes flow.
@@ 3.2.1 Linear stability analysis
where the constant $c$ can vary depending on the specific flow, but is ${\cal O}(1)$ (Goldman et al., 1967a; Ghalia et al., 2016), cf. Sec. 2.3. Near the fixed point of the Faxen field, the $\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}$ term in $W_{y}$ dominates the others ($\frac{\mathcal{B}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{3}),\frac{\mathcal{C}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{5}),\frac{\mathcal{D}_{large}}{\mathcal{A}_{large}}=\mathcal{O}(a_{p}^{4})$) and is the only one contributing to $\mathcal{O}(a_{p}^{3})$ in (18). This term is still of higher $a_{p}$ order than the Faxen correction so that, in the limit of large $\Delta$ and small $a_{p}$, the wall correction is perturbatively small.
@@ 3.3 Particle motion and manipulation in a clockwise vortex
Despite these caveats, the ability to concentrate particles on a limit cycle trajectory without inertial effects purely because of the background flow geometry is of fundamental interest. It is encouraging that the location of this limit cycle for small particles can be determined entirely within the large $\Delta$ approximation, i.e., without the intricate details of near-wall corrections or lubrication limits. This gives confidence in not only the qualitative but also the quantitative description of the phenomenon: when placed in certain bounded vortical Stokes flows, small spherical particles, even when neutrally buoyant, will eventually accumulate on well-defined closed trajectories.
@@ 3.4 Particle motion and manipulation in a counterclockwise vortex
Figure 6(b) exemplifies a particle trajectory that approaches very close to the wall, with a nearly normal initial approach and a subsequent piece of trajectory nearly parallel to the wall. Here, $\Delta\ll 1$ characterizes the trajectory and this is the only scenario in the present work where the variable expansion method for particle normal velocity (15) is necessary. The semi-logarithmic plot vs. time in Fig. 6(c) shows that the ”near-parallel” portion contains a stretch of approximately exponential decay of the (relative) gap $\Delta$. Indeed, this behavior follows from a series to leading order in small $\Delta$ of (15), i.e., the wall-expansion limit of $y_{E}\to-1$. Here, (14) simplifies to
@@
Our single-particle formalism implictily assumes the limit of dilute particle concentration in a microfluidics application. Including effects of particle-particle interaction is a desirable future extension of this approach in order to assess non-inertial effects in particle-laden flows important to practical applications (Guha, 2008).
@@ 4 Conclusions
We have quantitatively shown that force-free particles placed in an internal Stokes flow can be systematically displaced in a variety of ways through purely hydrodynamic interactions with the enclosing walls. Since a particle cannot contact a wall in finite time, a typical trajectory has a portion of approaching the wall and a portion of receding, governed by the background flow and the wall corrections derived from it. If the approaching and receding parts of the flow are symmetric, displacement effects off a streamline will cancel out. Thus, a Stokes-flow-inducing net particle displacement must break wall-parallel symmetry.
For practical use, vortical flows are advantageous, where the effects of many approaching/receding events can accumulate. However, if the vortex is symmetrically confined between walls, the net displacement effects induced by both walls will again cancel, and the particle trajectory must close. For net cumulative displacement, the vortex must, therefore, break wall-normal symmetry as well, for example by being confined between no-slip and no-stress boundaries. A net-transport internal Stokes flow thus breaks symmetries in both directions by particle-wall interactions, which act as perturbations on the Faxen flow trajectories. This is in agreement with recent experimental results that record permanent displacements of fibers passing by obstacles of symmetry-broken shape (Li et al., 2024). For such non-spherical particles (whether rigid or elastic), the effect of a single particle-boundary encounter could be much enhanced, as the wall effects affect different parts of the particle differently, and the overall torque balance will influence the resulting displacement. The study of this case will be the subject of future work.
@@
Let us estimate the rate of particle drift across streamlines in dimensional terms. Taking the length scale (channel half-width) as $H=50\mu$m, we want to ensure that the channel Reynolds number $Re=HU/\nu$ for aqueous solutions ($\nu\sim 10^{-6}$m2/s) is less than $10^{-2}$ along the particle trajectories, even in regions of the highest speed $U$. Given this constraint and assuming $a_{p}=5\mu$m, the time scale $\tau$ from (24) is $\approx 57$s so that a significant radial displacement is expected on a time scale of minutes. For larger particles, this time scale decreases rapidly ($\propto a_{p}^{-3}$), and the process can also be sped up by using a solution of higher viscosity.
The Moffatt eddy flow discussed here is convenient because of its closed analytical form. However, it would be extremely difficult to set up experimentally due to the exponential decay of the flow speed with $x$ and the need to drive the flow far away from the field of view. The exponential dependence of speed on coordinates also makes the Reynolds number constraint for Stokes flow extremely stringent. Nevertheless, the principles of displacement and the scaling of the wall effects are not specific to this flow and will be robust in any symmetry-breaking vortical Stokes flow, and the modeling equations for the crucial wall-normal displacement such as (11) and (15) are valid for arbitrary background flow.
@@
It is instructive to compare the scaling of the migration velocity of small particles far from the wall ($\Delta\gg 1$) with that of approaches using particle inertia. In a number of approaches quantifying the effect of steady inertia (Ho & Leal, 1974; Schonberg & Hinch, 1989; Asmolov & McLaughlin, 1999; Di Carlo et al., 2007; Hood et al., 2015) the migration speed for force-free particles is proportional to $a_{p}^{3}$, the same scaling as our result (18). In approaches using oscillatory inertia (Agarwal et al., 2021, 2024), the scaling varies from $a_{p}^{3}$ for small Stokes number to $a_{p}^{4}$ for large Stokes number. Thus, the strictly non-inertial effects described in the present work show scaling equal to or even more favorable than inertial techniques for small particles. It should be noted that at low (non-zero) $Re$ the walls of a microfluidic device will almost always be close enough to the particle to enable modeling as in Ho & Leal (1974) or Hood et al. (2015), i.e., the effects of inertia at large distances are not present. This is why Saffman lift (Saffman, 1965), an unbounded-flow effect, should not be directly compared with these results.